To determine the acceleration due to gravity applying linear least square
regression method by using a simple pendulum.
1
Expert's answer
2021-10-06T16:14:02-0400

The period T of a simple pendulum is represented by the formula, "T=2\pi \sqrt{L/g}"

Now, square both sides

"T^2=\frac{4\pi^2 L^2}{g}"

"g=\frac{4\pi^2 L^2}{T^2}"